# Balance-Based Cryptography: Physically Computing Any Boolean Function

### v1

@@ Abstract
Secure multi-party computation is an area in cryptography which studies how multiple parties can compare their private information without revealing it. Besides digital protocols, many physical protocols for secure multi-party computation using portable objects found in everyday life have also been developed. The vast majority of them use cards as the main tools. In this paper, we introduce the use of a balance scale and coins as new physical tools for secure multi-party computation. In particular, we develop four protocols that can securely compute any $n$-variable Boolean function using a balance scale and coins.
@@ 1 Introduction
Secure multi-party computation (MPC) is one of the most actively studied areas in cryptography, which investigates how multiple parties can compare their private information without revealing it.
One of the most classical problems in MPC involves the following setting. Alice and Bob want to know whether they both like each other. However, no one wants to confess first due to fear of embarrassment if getting rejected. They need a protocol that only distinguishes the two cases where they both like each other and otherwise, without leaking any other information. Theoretically, this setting is equivalent to computing a logical AND function of two input bits, one from each player.
@@
Instead of digital protocols, many researchers have developed physical protocols for MPC using portable objects found in everyday life such as cards, coins, and envelopes. These protocols have the benefit that they do not require computers and also allow external observers to verify that all parties truthfully execute them (which is often a challenging task for digital protocols). In addition, they are easier to understand and verify the correctness and security, even for non-experts in cryptography, and thus can be used for didactic purpose.
The vast majority of existing physical protocols use cards as the main tools, thus this area of research is often called card-based cryptography. During the past decade, card-based cryptography has rapidly gained interest from researchers and has been extensively studied, with dozens of papers published [6, 7, 17].
@@ 1.1 Related Work
Research in card-based cryptography began in 1989 when the five-card trick protocol was introduced by den Boer [4]. The five-card trick can compute the AND function of two input bits from two players using five cards: three identical black cards and two identical red cards.
Mizuki and Sone [12] later improved the AND protocol to accommodate $n$ inputs. Apart from the AND function, card-based protocols to compute other Boolean functions have also been proposed, including the XOR function [12], the majority function [16, 20], and the equality function [18]. Nishida et al. [15] proved that any $n$-variable Boolean function can be computed using $62n+$ 6 cards, and any such function that is symmetric can be computed using $22n+$ 2 cards.
Besides cards, other physical objects have been used in crptographic protocols as well, such as coins [8, 9], combination locks [11], envelopes [13], balls and bags [10], printed transparencies [3], and PEZ candy dispensers [1, 2, 14].
Very recently, Kaneko et al. [5] were the first ones to use a balance scale and coins in a cryptographic protocol. [^1] While earlier protocols [8, 9] also used coins, they used only head or tail faces to encode inputs without considering the weight of the coins. Their protocols, however, are not designed to compute functions in the MPC setting, but are zero-knowledge proof protocols to verify solutions of logic puzzles like Sudoku.
@@ 1.2 Our Contribution
In this paper, we introduce the use of a balance scale and coins as physical tools to securely compute Boolean functions, providing a new physical model for the MPC setting. Note that coins can be replaced by any small objects with significant weights like balls or marbles, but we use coins to preserve the setting proposed by Kaneko et al. [5].
@@
Unlike the protocols of Kaneko et al. [5] which use same-size coins with $n$ different weights, our protocols only use coins with two different weights, removing a practical challenge to implement them in real world.
@@
Protocol 3 (§5) symmetric $n\lceil\frac{n}{2}+2\rceil$ $\lceil\frac{n}{2}+1\rceil$ $\lceil\frac{n}{2}\rceil$ Uses a pen.
Protocol 4 (§6) any $n(2^{n}+1)$ $2^{n-1}$ $2^{n-1}$
@@ 2.1 Balance Scale
We use a double-pan balance scale in our protocols (see Fig. 1). When placing objects on both pans, the scale will tilt towards the heavier side at a fixed angle, or will stay put if the objects on both pans have equal weight.
@@ 2.2 Coins
Two types of coin, called heavy coins and light coins, are used in our protocols (see Fig. 1). All coins have the same size and color. Each heavy coin and light coin has weight $w$ and $w-\delta$, respectively. We assume that $\delta$ is much larger than $\varepsilon$, so the balance scale can detect the difference between heavy and light coins.
@@ 2.4 Bags and Shuffle Operation
Bags used in our protocols are small opaque bags that can be closed at the top with the same size, color, and weight (see Fig. 1). Each bag can contain multiple coins and can be placed on a balance scale.
@@ 3 Computing AND Function
3.1 Two-Variable
First, we propose the following simple protocol for one of the most fundamental functions: the two-variable AND function. In a setting with two players, each having a bit $x_{1}$ and $x_{2}$, respectively, this protocol securely computes $x_{1}\wedge x_{2}$.
@@ 3.1 Two-Variable
We now have two coins placed on one side of the scale. Then, the players jointly place two heavy coins on the other side of the scale. If the weights of the two sides are equal, return 1; otherwise, return 0.
This protocol is correct because the only case where the weights are equal is when both players’ coins are heavy coins (and the players’ coins are lighter in all other cases), which occurs when both players’ bits are 1s. Moreover, each player learns no information other than whether the two coins combined have weight $2w$ or less than $2w$. Therefore, the protocol is correct and secure. This protocol uses six coins (two given to each player plus two additional heavy coins) and uses one comparison.
3.2 $n$-Variable
Consider a setting where there are $n$ players, with each $i$-th player having a bit $x_{i}$. We can easily generalize the two-variable AND protocol to compute the $n$-variable AND function $x_{1}\wedge x_{2}\wedge...\wedge x_{n}$.
As in the case of two players, each player is given a heavy coin and a light coin, and then chooses the heavy (resp. light) coin if the player’s bit is 1 (resp. 0) to place on the same side of a balance scale.
The players then place $n$ heavy coins on the other side of the scale. If the weights of the two sides are equal, return 1; otherwise, return 0.
@@ 4 Computing Threshold Function
The players then place a custom weight $W=(k-1)w+(n-k+1)(w-\delta)+\frac{\delta}{2}$ on the other side of the scale. [^2] Alternatively, the weight $W$ can be obtained by placing $k-1$ heavy coins, $n-k+1$ light coins, and a custom weight of $\frac{\delta}{2}$. If the $n$ coins combined are heavier than $W$, return 1; otherwise, return 0.
This protocol is correct because the $n$ coins are heavier than $W$ if and only if at least $k$ of them are heavy coins (and the $n$ coins are lighter in all other cases). This occurs when at least $k$ of the players’ bits are 1s, i.e. $\sum_{i=1}^{n}x_{i}\geq k$, which is the exact condition for $T_{k}(x_{1},x_{2},...,x_{n})=1$. Moreover, each player learns no information other than whether the $n$ coins are heavier or lighter than $W$. Therefore, the protocol is correct and secure.
@@ A Boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ is called symmetric if
for any $x_{1},x_{2},...,x_{n}\in\{0,1\}$ and any permutation $(\sigma(1),\sigma(2),...,\sigma(n))$ of $(1,2,...,n)$. Note that for any such function $f$, the value of $f(x_{1},x_{2},...,x_{n})$ only depends on the number of inputs being 1s, i.e. the integer sum $\sum_{i=1}^{n}x_{i}$.
@@
To compute $S_{X}^{n}$, we prepare $|X|$ bags, each one corresponding to each element of $X$. A bag corresponding to $k\in X$ contains $k$ heavy coins and $n-k$ light coins.
@@ 5.1 Proofs of Correctness and Security
The protocol returns 1 when there is a non-special bag with weight equal to the special bag (and there can be at most one such bag). Suppose this non-special bag corresponds to an element $k\in X$ (containing $k$ heavy coins and $n-k$ light coins). This case occurs if and only if exactly $k$ of the $n$ players’ coins are heavy coins, which is equivalent to $\sum_{i=1}^{n}x_{i}=k$. Therefore, the protocol returns 1 if and only if $\sum_{i=1}^{n}x_{i}\in X$, which is the exact condition for $S_{X}^{n}(x_{1},x_{2},...,x_{n})=1$.
Consider the process of comparing a special bag and a non-special bag in each iteration. Due to the shuffles both before and after weighing, no player can distinguish which side of the scale is a special bag. Therefore, all information each player learns is whether or not both bag have equal weight.
Also, due to the shuffle of all $|X|$ non-special bags into a uniformly random permutation, no player can learn about which non-special bag correspond to which element of $X$. Therefore, all information each player learns is whether or not there is a non-special bag with weight equal to the special bag. Therefore, the protocol is correct and secure.
@@ 5.2 Optimization
In the protocol, we need $|X|+1$ bags to compute $S_{X}^{n}$. However, if $|X|>\frac{n+1}{2}$, we can instead compute the function $S_{\{0,1,...,n\}-X}^{n}$ and then negate the output. In such cases, we have |$|\{0,1,...,n\}-X|=n+1-|X|<\frac{n+1}{2}$. Therefore, the number of required bags can be reduced to at most $\lfloor\frac{n+1}{2}\rfloor+1=\lceil\frac{n}{2}\rceil+1$.
@@ 6 Computing Any Boolean Function
Finally, we propose the following protocol to compute an arbitrary $n$-variable Boolean function among $n$ players, where each $i$-th player has a bit $x_{i}$.
@@
$=\bigvee_{(b_{1},b_{2},...,b_{n})\in\{0,1\}^{n}}(\bar{x}_{1}\oplus b_{1})(\bar{x}_{2}\oplus b_{2})...(\bar{x}_{n}\oplus b_{n})f(b_{1},b_{2},...,b_{n}),$
where $\vee$ and $\oplus$ are Boolean OR and XOR operations, respectively. Observe that we only need to consider the terms where $f(b_{1},b_{2},...,b_{n})=1$, since the rest of the terms are all zeroes. The output $f(x_{1},x_{2},...,x_{n})$ is 1 if and only if is at least one such term is 1.
@@
For each $i$-th player and each bag corresponding to $(b_{1},b_{2},...,b_{n})$, the player puts in that bag a heavy coin if $x_{i}=b_{i}$, and a light coin if $x_{i}\neq b_{i}$.
@@ 6.2 Proofs of Correctness and Security
The protocol returns 1 when there is a bag corresponding to $(b_{1},b_{2},...,b_{n})\in B_{f}$ containing $n$ heavy coins, which occurs when $x_{i}=b_{i}$ for every $i=1,2,...,n$ (and there can be at most one such bag). Therefore, the protocol returns 1 if and only if $(x_{1},x_{2},...,x_{n})\in B_{f}$, which is the exact condition for $f(x_{1},x_{2},...,x_{n})=1$.
@@ 6.3 Optimization
For $1\leq i\leq n$, let $p_{i}$ (resp. $q_{i}$) be the numbers of $(b_{1},b_{2},...,b_{n})\in B_{f}$ such that $b_{i}=0$ (resp. $b_{i}=1$). Observe that the $i$-th player has to use at most $\max(p_{i},q_{i})$ heavy coins and at most $\max(p_{i},q_{i})$ light coins (we have $\max(p_{i},q_{i})\leq|B_{f}|\leq 2^{n-1}$). As a result, we can give only $\max(p_{i},q_{i})$ heavy coins and $\max(p_{i},q_{i})$ light coins to each $i$-th player (instead of the maximum $2^{n-1}$ heavy coins and $2^{n-1}$ light coins). For instance, in the example in Section 6.1, each player receives only two heavy coins and two light coins (instead of four heavy coins and four light coins). The total number of required coins is at most $n(2^{n}+1)$ (at most $2^{n}$ given to each player plus $n$ additional heavy coins).
7 Future Work
We introduced the use of a balance scale and coins in the MPC setting by developing four balance-based protocols that securely compute $n$-variable Boolean functions. An interesting future work is to develop a protocol that can compute an arbitrary Boolean function with the polynomial numbers of coins and bags, or to prove the lower bound of those numbers.

### v2

@@ Abstract
Secure multi-party computation is an area in cryptography which studies how multiple parties can compare their private information without revealing it. Besides digital protocols, many unconventional protocols for secure multi-party computation using physical objects have also been developed. The vast majority of them use playing cards as the main tools. In 2024, Kaneko et al. introduced the use of a balance scale and coins in zero-knowledge proof protocols for pencil puzzles. In this paper, we extend the use of these tools to secure multi-party computation. In particular, we develop four protocols that can securely compute any $n$-variable Boolean function using a balance scale and coins.
@@ 1 Introduction
Secure multi-party computation (MPC) is one of the most actively studied areas in cryptography. It investigates how multiple parties can compare their private information without revealing it.
One of the best-known classical examples of MPC involves the following problem. Alice and Bob want to know whether they both like each other. However, no one wants to confess first due to fear of embarrassment of getting rejected. They need a protocol that only distinguishes the two cases where they both like each other and otherwise without leaking any other information. Theoretically, this setting is equivalent to computing a logical AND function of two input bits, one from each player.
@@
Instead of digital protocols, many researchers have developed unconventional protocols for MPC using physical objects found in everyday life such as cards, coins, and envelopes. These protocols have the benefit that they do not require computers and also allow external observers to verify that all parties truthfully execute them (which is often a challenging task for digital protocols). In addition, they are easier to understand and to verify the correctness and security, even for non-experts, and thus can be used for didactic purposes.
The vast majority of existing physical protocols use playing cards as the main tools, thus this area of research is often called card-based cryptography. During the past decade, card-based cryptography has rapidly gained interest and has been extensively studied, with dozens of papers published [7, 8, 19].
@@ 1.1 Related Work
Research in card-based cryptography began in 1989 when den Boer [4] introduced the five-card trick protocol to solve the classical Alice-and-Bob problem. This protocol can compute the AND function of two input bits from two players using five cards: three identical black cards and two identical red cards.
Mizuki and Sone [14] later improved the AND protocol to accommodate $n$ inputs. Apart from the AND function, card-based protocols to compute other Boolean functions have also been proposed, including the XOR function [14], the majority function [18, 22], and the equality function [20]. Nishida et al. [17] proved that any $n$-variable Boolean function can be computed using $2n+6$ cards, and any such function that is symmetric can be computed using $2n+2$ cards.
Besides cards, other physical objects such as coins [9, 10], envelopes [15], combination locks [12], balls and bags [11], printed transparencies [3], and PEZ candy dispensers [1, 2, 16] have also been used in cryptographic protocols.
In 2024, Kaneko et al. [5] were the first ones to use a balance scale and coins in a cryptographic protocol. [^1] While some earlier protocols [9, 10] also used coins, only faces (head or tail) were used to encode inputs without considering the weight of the coins. Their protocols, however, were not designed to compute functions in the MPC setting, but are zero-knowledge proof protocols to verify solutions of pencil puzzles like Sudoku. Very recently, Kaneko et al. [6] also proposed a voting protocol using the same tools, which can compute a specific function in the MPC setting.
@@ 1.2 Our Contribution
In this paper, we extend the use of a balance scale and coins as physical tools to securely compute arbitrary Boolean functions, providing a new physical model for a generic MPC setting. Note that coins can be replaced by any small objects with significant weights like balls or marbles, but we use coins to preserve the setting proposed by Kaneko et al. [5].
@@
Our protocols are also practical. Unlike the two existing balance-based protocols [5, 6] which use same-size coins with $n$ different weights, our protocols only use coins with two different weights, removing a practical challenge to implement them in real world.
@@
Protocol 3 (§5) symmetric at most $n\lceil\frac{n}{2}+2\rceil$ at most $\lceil\frac{n}{2}+1\rceil$ at most $\lceil\frac{n}{2}\rceil$ Uses a pen.
Protocol 4 (§6) any at most $n(2^{n}+1)$ at most $2^{n-1}$ at most $2^{n-1}$
@@ 2.1 Balance Scale
Our protocols use a double-pan balance scale (see Fig. 1). When placing objects on both pans, the scale will tilt towards the heavier side at a constant angle (regardless of weight difference), or will stay put if the objects on both pans have equal weight.
@@ 2.2 Coins
Two types of coin, called heavy coins and light coins, are used in our protocols. All coins have the same size and color (see Fig. 1). Each heavy coin weighs $w$, while each light coin weighs $w-\delta$. We assume that $\delta$ is much larger than $\varepsilon$, so the balance scale can detect the difference between heavy and light coins.
@@ 2.4 Bags and Shuffle Operation
Bags used in our protocols are small opaque bags with the same size, color, and weight, and can be closed at the top (see Fig. 1). Each bag can contain multiple coins and can be placed on a balance scale.
@@ 3 Computing AND Function
First, we propose the following protocol for one of the most fundamental functions: the AND function. In a setting with $n$ players, where the $i$-th player has a bit $x_{i}$, this protocol securely computes $x_{1}\wedge x_{2}\wedge...\wedge x_{n}$.
@@
We now have $n$ coins placed on one side of the scale. Then, the players jointly place $n$ heavy coins on the other side of the scale. If the weights of the two sides are equal, return 1; otherwise, return 0.
@@ 4 Computing Threshold Function
The players then place a custom weight $W=(k-1)w+(n-k+1)(w-\delta)+\frac{\delta}{2}$ on the other side of the scale. [^2] Alternatively, the weight $W$ can be obtained by placing k1k-1 heavy coins, nk+1n-k+1 light coins, and a custom weight of $\frac{\delta}{2}$. If the $n$ coins combined are heavier than $W$, return 1; otherwise, return 0.
Observe that $W$ is heavier than k1k-1 heavy coins plus nk+1n-k+1 light coins, but is lighter than $k$ heavy coins plus nkn-k light coins. As a result, the $n$ coins are heavier than $W$ if and only if at least $k$ of them are heavy coins (and the $n$ coins are lighter in all other cases). This occurs when at least $k$ of the players’ bits are 1s, i.e. $\sum_{i=1}^{n}x_{i}\geq k$, which is the exact condition for $T_{k}(x_{1},x_{2},...,x_{n})=1$. Moreover, each player learns no information other than whether the $n$ coins are heavier or lighter than $W$. Therefore, the protocol is correct and secure.
@@ A Boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ is called symmetric if
for any $x_{1},x_{2},...,x_{n}\in\{0,1\}$ and any permutation $\sigma:\{1,2,...,n\}\rightarrow\{1,2,...,n\}$. Note that for any such function $f$, the value of $f(x_{1},x_{2},...,x_{n})$ only depends on the number of inputs being 1s, i.e. the integer sum $\sum_{i=1}^{n}x_{i}$.
@@
To compute $S_{X}^{n}$, we prepare $|X|$ bags, each one corresponding to each element of $X$. A bag corresponding to $k\in X$ contains $k$ heavy coins and nkn-k light coins.
@@ 5.1 Proofs of Correctness and Security
The protocol returns 1 when there is a non-special bag with weight equal to the special bag (and there can be at most one such bag). Suppose this non-special bag corresponds to an element $k\in X$ (containing $k$ heavy coins and nkn-k light coins). This case occurs if and only if exactly $k$ of the $n$ players’ coins are heavy coins, which is equivalent to $\sum_{i=1}^{n}x_{i}=k$. Therefore, the protocol returns 1 if and only if $\sum_{i=1}^{n}x_{i}\in X$, which is the exact condition for $S_{X}^{n}(x_{1},x_{2},...,x_{n})=1$.
Consider the process of comparing a special bag and a non-special bag in each iteration. Due to the shuffles both before and after weighing, no player can distinguish which side of the scale is a special bag. Therefore, all information each player learns is whether or not both bags have equal weight.
Also, due to the shuffle of all $|X|$ non-special bags into a uniformly random permutation, no player can learn about which non-special bag corresponds to which element of $X$. Therefore, all information each player learns is whether or not there is a non-special bag with weight equal to the special bag. Therefore, the protocol is correct and secure.
@@ 5.2 Optimization
In the protocol, we need $|X|+1$ bags to compute $S_{X}^{n}$. However, if $|X|>\frac{n+1}{2}$, we can instead compute the function $S_{\{0,1,...,n\}-X}^{n}$ and then negate the output. In such cases, we have $|\{0,1,...,n\}-X|=n+1-|X|<\frac{n+1}{2}$. Therefore, the number of required bags can be reduced to at most $\lfloor\frac{n+1}{2}\rfloor+1=\lceil\frac{n}{2}\rceil+1$.
@@ 6 Computing Any Boolean Function
Finally, we propose the following protocol to compute an arbitrary $n$-variable Boolean function among $n$ players, where the $i$-th player has a bit $x_{i}$.
@@
$=\bigvee_{(b_{1},b_{2},...,b_{n})\in\{0,1\}^{n}}(x_{1}\leftrightarrow b_{1})(x_{2}\leftrightarrow b_{2})...(x_{n}\leftrightarrow b_{n})f(b_{1},b_{2},...,b_{n}),$
where $\vee$ and $\leftrightarrow$ are Boolean OR and XNOR (i.e. if and only if) operations, respectively. Observe that we only need to consider the terms where $f(b_{1},b_{2},...,b_{n})=1$, since the rest of the terms are all zeroes. The output $f(x_{1},x_{2},...,x_{n})$ is 1 if and only if at least one such term is 1.
@@
For each bag corresponding to $(b_{1},b_{2},...,b_{n})$, the $i$-th player puts into that bag a heavy coin if $x_{i}=b_{i}$, and a light coin if $x_{i}\neq b_{i}$. [^3] We assume a semi-honest model [13], where each player puts a coin into every bag consistently according to the rule and does not try to sabotage or hack the protocol.
@@
Note that for a special case where $|B_{f}|=1$, a bag is not necessary; the players can directly place coins on the scale similarly to the AND protocol.
@@ 6.2 Proofs of Correctness and Security
The protocol returns 1 when there is a bag corresponding to $(b_{1},b_{2},...,b_{n})\in B_{f}$ containing $n$ heavy coins, which occurs when $x_{i}=b_{i}$ for every i=1,2,,ni=1,2,...,n (and there can be at most one such bag). Therefore, the protocol returns 1 if and only if $(x_{1},x_{2},...,x_{n})\in B_{f}$, which is the exact condition for $f(x_{1},x_{2},...,x_{n})=1$.
@@ 6.3 Optimization
For $1\leq i\leq n$, let $p_{i}$ (resp. $q_{i}$) be the numbers of $(b_{1},b_{2},...,b_{n})\in B_{f}$ such that $b_{i}=0$ (resp. $b_{i}=1$). Observe that the $i$-th player has to use at most $\max(p_{i},q_{i})$ heavy coins and at most $\max(p_{i},q_{i})$ light coins (we have $\max(p_{i},q_{i})\leq|B_{f}|\leq 2^{n-1}$). As a result, we can give only $\max(p_{i},q_{i})$ heavy coins and $\max(p_{i},q_{i})$ light coins to the $i$-th player (instead of the maximum $2^{n-1}$ heavy coins and $2^{n-1}$ light coins). For instance, in the example in Section 6.1, each player receives only two heavy coins and two light coins (instead of four heavy coins and four light coins). The total number of required coins is at most $n(2^{n}+1)$ (at most $2^{n}$ given to each player plus $n$ additional heavy coins).
7 Conclusion
We developed four balance-based protocols that can compute arbitrary $n$-variable Boolean functions, and proved the correctness and security of these protocols. By doing so, we extended the use of a balance scale and coins to a generic MPC setting, implying that they have computational power equal to cards, PEZ candy dispensers, and other existing physical tools.